OpenStudy (anonymous):

a salesperson at a car dealership has a base salary of $25,000 per year and earns a 5% commision on total sales. how much must the salesperson sell to earn $50,000 in one year?

OpenStudy (anonymous):

Therefore the salesperson must sell $500 000 to earn $50 000 in one year

OpenStudy (anonymous):

50000 = 25000 + 0.05x Now, we solve for x: 50000 - 25000 = 0.05x 25000 = 0.05x 500 000 = x

OpenStudy (anonymous):

-2(4-3x)+7=6(x+1)